Please pleeease help me. I have been struggling for 3 days to find the answer.

I NEED ANSWER FOR Budgeted Manufacturing Cost per Unit account and Budgeted Income statement and balance sheet. Please show calculations if possible. Thank you Current assets as of December 31 (prior year):  

     Cash $4,650.00, Accounts receivable, net $57,600.00, Inventory   $15,600.00; Property, plant, and equipment, net $121,500.00; Accounts payable   $42,800.00;   Capital stock $124,500.00; Retained earnings   $22,800.00

A:            Actual Sales in December were$72,000. Selling price per unit projected to remain stable at $12 per unit throughout the budget period. Sales for the first five months of the upcoming year are budgeted to be as follow: January    $ 104,400    , February $ 108,000 , March $ 112,800 , April $ 109,200, May $ 105,600

B:            Sales are 20% cash and 80% credit. All credit sales are collected in the months following the sale.                                                 

C:            the company has a policy that states that each months ending inventory of finished goods should be 10% of the following month's sales (in units)                             

D:            Of each month's direct material purchases, 20% are paid for in the month of purchase, while the remainder is paid for in the month following purchase. 3kg of direct material is needed at $2.00 per kg. Ending inventory of direct materials should be 30% of next month's production needs

E;             Monthly manufacturing conversion costs Factory rent $ 4,500.00, Other fixed manufacturing expenses $ 2,800.00, Variable manufacturing overhead ($1.10 per unit) $ 1.10, No depreciation is included in these figures. All expenses are paid in the month in which they are incurred

F:    Computer equipment for administrative offices will be purchased in the upcoming quarter incurred, In January, Osborne Manufacturing will purchase January purchase - equipment for $6000 (cash). February cash expenditure will $ 12,800.00, March cash expenditure $15,600.00

g:             Operating expenses are budgeted to be $1.30 per unit sold     $ 1.30 , Plus fixed operating expenses of 1800 per month      $ 1,800.00. All operating expenses are paid in the month in which they are incurred.

H:      Depreciation on the building and equipment for the general administrative offices is budgeted to $4600 for the entire quarter, which includes depreciation on new acquisitions. Depreciation on the building and equipment budget $4,600.00

I:              The Company has a policy that the ending cash balance in each month must be at least $4200, The Company has a line of credit with a local bank. It can borrow increments of $1000 at the beginning of each month up to a total outstanding loan balance of $130,000. The interest rate on these loans is 2% per month simple interest. The company pays down on the line of credit balance if it has excess funds at the end of the quarter. The company also pays the accumulated interest at the end of the quarter on the funds borrowed during the      J:      The company's income tax is projected to be 30% of operating income less interest expense. The company pays $10800 cash at the end of February in estimated taxes   
Answer #1

Cash collection budget

January

February

March

Quarter

Cash sales

20880

21600

22560

65040

Credit sales

57600

83520

86400

227520

Total cash collections

78480

105120

108960

292560

January

Cash = 104400*20% =20880

Credit = 72000*80% =5700

February

Cash = 108000*20% =21600

Credit = 104400*80% =83520

March

Cash = 112800*20% =22560

Credit = 108000*80% =86400

Production budget

January

February

March

Quarter

Unit sales

8700

9000

9400

27100

Plus: Desired ending inventory (next month’s unit sales *10%)

900

940

910

910

Total needed

9600

9940

10310

28010

Less: Beginning inventory (current month’s unit sales *10%)

870

900

940

870

Number of units to produce

8730

9040

9370

27140

January

Unit sales = 104400/12 = 8700

February

Unit sales = 108000/12 = 9000

March

Unit sales = 112800/12 = 9400

April

Unit sales = 109200/12 = 9100

Direct materials budget

January

February

March

Quarter

Units to be produced

8730

9040

9370

27140

Multiply by: Quantity (pounds) of DM needed per unit

3

3

3

3

Quantity (pounds) needed for production

26190

27120

28110

81420

Plus: Desired ending inventory of DM (next month’s quantity needed for production *30%)

8136

8433

8163

8163

Total quantity (pounds) needed

34326

35553

36273

89583

Less: Beginning inventory of DM

7857

8136

8433

7857

Quantity (pounds) to purchase

26469

27417

27840

81726

Multiply by: cost per pound

2

2

2

2

Total cost of DM purchases

52938

54834

55680

163452

Units to be produced in April = 9100+(105600/12*10%)-910 = 9070

Quantity (pounds) needed for production in April = 9070*3 =27210

Cash payments for direct materials purchase

January

February

March

Quarter

December purchases (From Accounts Payable)

42800

42800

January purchases

10588

42350

52938

February purchases

10967

43867

54834

March purchases

11136

11136

Total payments

53388

53317

55003

161708

January

52938*20% = 10588

52938*80% = 42350

February

54834*20% = 10967

54834*80% = 43867

March

55680*20% = 11136

Cash payments for manufacturing overhead budget

January

February

March

Quarter

Rent (fixed)

4500

4500

4500

13500

Other MOH (fixed)

2800

2800

2800

8400

Variable manufacturing overhead costs (unit to be produced *1.10)

9603

9944

10307

29854

Total disbursements

16903

17244

17607

51754

Cash payments for operating expense budget

January

February

March

Quarter

Variable operating expenses (unit sales *1.30)

11310

11700

12220

35230

Fixed operating expenses

1800

1800

1800

5400

Total payments for operating expenses

13110

13500

14020

40630

Combine cash budget

January

February

March

Quarter

Cash balance, beginning

4650

4729

5188

4650

Plus: cash collections

78480

105120

108960

292560

Total cash available

83130

109849

114148

297210

Less cash payments:

DM purchases

53388

53317

55003

161708

MOH costs

16903

17244

17607

51754

Operating expenses

13110

13500

14020

40630

Tax payment

10800

10800

Equipment purchases

6000

12800

15600

34400

Total cash payments

89401

107661

102230

299292

Ending cash before financing

(6271)

2188

11918

(2082)

Financing:

Borrowings

11000

3000

14000

Repayments

(6000)

Interest payments

(780)

(6000)

Total financing

11000

3000

(6780)

7220

Cash balance, ending

4729

5188

5138

5138

Budgeted Manufacturing Cost per Unit

Direct materials cost per unit

$6.00

Variable manufacturing costs per unit

$1.10

Fixed manufacturing overhead per unit

$0.80

Cost of manufacturing each unit

$7.90

(13500+8400)/ 27140 = $0.80

Budgeted Income Statement

Sales (104400+108000+112800)

325200

Less: Cost of goods sold (27100*7.90)

(214090)

Gross profit

111110

Less: Operating expenses

(40630)

Less: Depreciation expense

(4600)

Operating income

65880

Less: interest expense

(780)

Less: income tax expense @ 30% (65880-780)*30%)

(19530)

Net income

$45570